OpenStudy (anonymous):

A ladder, 10 ft long, rests against a wall. If the bottom of the ladder slides away from the wall at 2 ft/sec, how fast is the angle between the top of the ladder and the wall changing when the angle is 45 degrees?

OpenStudy (baru):



OpenStudy (baru):

can you write \(\theta\) in terms of x?

OpenStudy (anonymous):

\[\sin \Theta=\frac{ x }{ 10 }\] \[x=\sin \Theta \times10\]

OpenStudy (baru):

now differentiate both sides with respect to 't'

OpenStudy (anonymous):

\[\frac{ dx }{ dt }=\cos \theta \times \frac{ d \theta }{ dt }(10)\]

OpenStudy (baru):

read the question you are given,dx/dt and the angle theta=45 substitute

OpenStudy (anonymous):

Oh ok! \[100=\cos(45)\times10(\frac{ d \theta }{ dt }\]

OpenStudy (anonymous):

So all that's left is solve for the derivative of theta?

OpenStudy (baru):

dx/dt=2 ft/sec re-check, how did u get 100?

OpenStudy (anonymous):

Oh, my bad. I was looking at the wrong value. Saw the 10 ft in the problem and for some reason wrote 100

OpenStudy (baru):

cos(45)=\(\frac{1}{\sqrt{2}}\)

OpenStudy (baru):

substitute and find \(d\theta/dt\)

OpenStudy (anonymous):

Ok, thank you for the help!

OpenStudy (baru):

:)